Marlow Company purchased a point of sale system on January 1 for $7100. The system has a useful life for 5 years and a salvage value of $1250. What would be the depreciation expense for the second year of its useful life using the double-declining-balanced method?

A $2340 B $2840 C $1704 D $1640 E $19,550
Answer #1

Option (C) is correct

Under double declining balance method, depreciation is calculated by the following formula:

Depreciation = 2 * 1 / N * (Cost - Accumulated depreciation)

where, n is the no. of years or useful life and accumulated depreciation is depreciation charged till date

For first year, accumulated depreciation will be zero.

So, depreciation for first year will be:

Depreciation = 2 * 1 / 5 * ($7100 - $0)

Depreciation = 2 * 1 / 5 * $7100 = $2840

Now, in second year, accumulated depreciation will be $2840 (i.e. depreciation charged till date is depreciation for first year)

Depreciation for second year will be:

Depreciation = 2 * 1 / 5 * ($7100 - $2840) = 2 * 1 / 5 * $4260 = $1704